OpenStudy (anonymous):

Given the function f(x) = 3(x+2) − 4, solve for the inverse function when x = 2.

OpenStudy (trojanpoem):

Simplify : f(x) = 3x + 2 -4 f(x) = 3x - 2 Let f(x) = y y = 3x - 2 y + 2 = 3x x = (y+2)/3 swap x, y y = (x+2)/3 (That's the inverse) At x = 2 y = (2+2)/3 = 4/3

OpenStudy (trojanpoem):

Forgot to mention that the function isn't one to one f(-x) = -3x -2 (not equal to ) f(x)